OpenStudy (sleepyhead314):

Let f be the function defined by f(x)= 5cos(pix) - e^(x^2 -4). If g(x)= inverse of f(x), find the equation of the line tangent to the graph if g at point where g(x)=2 I would just like to check that I understood how to do this properly... derivative of f(x) = -5pisin(pix) -2xe^(x^2 -4) Plug in 2 for x -> -4 slopes of inverse functions are reciprocals so slope at g(x)=2 would be -1/4 Then plugging in 2 for x into f(x) because g(x)=2 means f(2)=x So gets point (4,2) in terms of g So I get y= -1/4 (x-4) +2 as the tangent line? :3

OpenStudy (sleepyhead314):

@whpalmer4 @sourwing @Luigi0210 @johnweldon1993 help check? :3

OpenStudy (sleepyhead314):

@agent0smith @campbell_st @RadEn could you pwease help me check? :3

OpenStudy (sleepyhead314):

@zepdrix @robtobey @sourwing

OpenStudy (sleepyhead314):

hmmm? I got f'(x) then said slopes of inverse functions are reciprocals

OpenStudy (sleepyhead314):

I know that f^-1(x) doesnt equal f'(x) I got f'(x) cuz g'(x) = 1/f'(g(x))

OpenStudy (anonymous):

you're basically calculating 1/f'(2)

OpenStudy (sleepyhead314):

yes, found f'(x) plugged in x=2 then reciprocal

OpenStudy (sleepyhead314):

don't have to do anything by hand, calculator question

OpenStudy (anonymous):

You did it right

OpenStudy (sleepyhead314):

ok thank you! :D <3

OpenStudy (anonymous):

except one minor mistake. it's at (-4,2). not at (4,2)

OpenStudy (anonymous):

y= -1/4 (x+4) +2 is the correct answer

OpenStudy (sleepyhead314):

why would it be -4 ? http://www.wolframalpha.com/input/?i=5cos%28pix%29+-+e%5E%28x%5E2+-4%29+for+x%3D2

OpenStudy (anonymous):

because (2,-4) is the point on f, then this means (-4,2) is the point on g(x)

OpenStudy (sleepyhead314):

but wolfram told me that point on f(2) = 4 ;-; http://www.wolframalpha.com/input/?i=5cos%28pix%29+-+e%5E%28x%5E2+-4%29+for+x%3D2

OpenStudy (anonymous):

ahh I see what's up now. I took your word for it. You said in the question that you got -4 when x = 2

OpenStudy (sleepyhead314):

for derivative :P f'(2) = -4

OpenStudy (sleepyhead314):

f(2) = 4 xD

OpenStudy (anonymous):

O.O ah crap. I should have read the question carefully :D.

OpenStudy (anonymous):

so (4,2) is on g. and 1/f'(f^-1(x)) = 1/f'(2) = 1/-4 = -1/4 yeah it's correct :DD

OpenStudy (sleepyhead314):

Thank you ^_^

OpenStudy (anonymous):

good job. You understand the concept ^.^b

OpenStudy (sleepyhead314):

I hope I do :P APs are in two weeks >.> O_O

OpenStudy (anonymous):

ah wait O.O

OpenStudy (sleepyhead314):

wut now? o.o stop scaring me ;-; xD

OpenStudy (anonymous):

Shouldn't we be calculating 1/f'(4) instead?

OpenStudy (anonymous):

oh no nmv XD

OpenStudy (sleepyhead314):

;-; meanie why you scare my like dat >,<

OpenStudy (anonymous):

nvm XD. This whole derivative of inverse thing can really drive you crazy XDD

OpenStudy (sleepyhead314):

lol ikr xD for some reason none of the APCalcBC people know it but apparently the APCalcAB kids are >.>

OpenStudy (sleepyhead314):

my grammer just failed me xD

OpenStudy (anonymous):

Grammar errors are acceptable in a math class :D

OpenStudy (sleepyhead314):

:D yrsh xD